# Open-world Machine Learning: A Review and New Outlooks

### v2

@@ 3.3 Analysis of Unknown Rejection
The ability of a classifier to reject unknowns has been verified as highly correlated with its closed-set classification accuracy [35]. As a result, when applying some training tricks to the classifier for better closed-set classification, e.g., longer training, RandAugment [80], learning rate warmup, and label smoothing [81], the classifier consequently performs better in OSR. The authors [35] further found that applying baseline methods like MSP [10] and MLS [44] on a well-trained closed-set classifier could achieve competitively with or even outperform state-of-the-art OSR methods. Cen et al. [82] further revealed that wrong-classified ID samples have similar uncertainty distribution to unknown samples, which limits the performance of OSR, and pre-training could enhance both the separation between ID and OOD samples and the separation between the correctly- and wrongly-classified ID samples.
Regarding the difficulty of unknown rejection, existing works predominantly resort to openness [31] $O^{\star}$ defined by the number of known and unknown classes, i.e., $O^{\star}=1-\sqrt{\frac{2\times K_{\text{ID}}}{K_{\text{ID}}+K_{\text{OOD}}}}$, where $K_{\text{ID}},K_{\text{OOD}}$ denote the number of known and unknown classes, respectively. The more unknown categories compared to the number of known categories, the more challenging the unknown rejection task becomes. Vaze et al. [35] proposed a more rational metric namely semantic similarity [83]. Open-set samples that share a greater number of visual features, such as attributes, color, and shape, with known training samples, are more difficult to reject.
4 Novel Class Discovery
@@
Novel class discovery [13, 84] is the second step in open-world machine learning, which aims to automatically discover novel categories from unlabeled data based on the model’s prior knowledge. Novel class discovery is an extension of unknown rejection, requiring models to not only reject unknown samples, but also further extrapolate to classify the rejected samples. Given a dataset $\mathcal{D}_{\text{train}}=\mathcal{D}_{l}\cup\mathcal{D}_{u}$ with two subsets, where $\mathcal{D}_{l}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N_{l}}\subset\mathcal{X}\times\mathcal{Y}_{l}$ is the labeled dataset and $\mathcal{D}_{u}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N_{u}}\subset\mathcal{X}\times\mathcal{Y}_{u}$ is the unlabeled data. It is worth noting that $\mathcal{Y}_{u}\neq\mathcal{Y}_{l}$, i.e., there are novel classes in the unlabeled data outside the old classes in $\mathcal{D}_{l}$. Let $\mathcal{C}_{\text{old}},\mathcal{C}_{\text{new}}$ denote the old and new classes in the labeled and unlabeled dataset, with the number of classes $K_{\text{old}}=|\mathcal{C}_{\text{old}}|,K_{\text{new}}=|\mathcal{C}_{\text{new}}|$, respectively. The objective is to cluster and discover $\mathcal{C}_{\text{new}}$ in $\mathcal{D}_{u}$ leveraging the knowledge learned on $\mathcal{D}_{l}$.
@@
In essence, novel class discovery is a clustering task on $\mathcal{D}_{u}$. Regarding its differences from unsupervised clustering [85, 86], the latter aims to cluster unlabeled data in a purely unsupervised manner without any prior knowledge of classification criterion, as a result, unsupervised clustering is not a fully learnable task. For example, in the absence of any external information, the model might face a dilemma, confusing whether to cluster red flowers and red cars together according to colors or to cluster red flowers and yellow flowers together according to species. From this perspective, some labeled data $\mathcal{D}_{l}$ is necessary and could provide some prior knowledge indicating what constitutes a class to clarify the clustering criterion and eliminate ambiguity. In a word, novel class discovery is a more reasonable and pragmatic task, and could be referred to as deep transfer clustering [87], models are supposed to transfer the knowledge learned from old and labeled classes to cluster novel categories in the unlabeled data.
@@
Key Challenges. Intrinsically, the semantic similarity between new and old categories determines the extent of knowledge transferable from old to new categories, thereby influencing the performance of category discovery. If the semantic similarity is limited, the labeled old class may degrade the performance novel class discovery [88]. Technically, novel category discovery could be referred to as open-world semi-supervised learning [89], where unlabeled samples contain new classes, which makes the conventional pseudo-labeling mechanism [90] not applicable due to the non-shared categories between labeled and unlabeled data. Moreover, in practical scenarios, the number of novel categories is always unknown, which makes the task challenging.
4.1 Novel Category Discovery
@@
Multi-stage Methods. Early works in NCD mainly transfer knowledge from $\mathcal{D}_{l}$ to $\mathcal{D}_{u}$ for category discovery in multi-steps. A pioneer work DTC [87] formalizes NCD as a deep transfer clustering task, and builds the method upon unsupervised clustering DEC [85]. DTC first trains a model on $D_{l}$ which is then utilized to extract features of $\mathcal{D}_{u}$ to initialize cluster centroids. Then DTC performs alternative optimization between the soft targets and model parameters together with cluster centroids. Later works predominantly resort to pseudo-labels to transfer knowledge from labeled samples from old classes to unlabeled samples from novel classes. However, due to the disjoint classes, one could not simply adopt the naive pseudo-labels [90] in semi-supervised learning [91]. To overcome this issue, several methods rely on pairwise pseudo-labels [92] indicating whether two samples are similar or not. Specifically, the model is firstly trained on $\mathcal{D}_{l}$ to assign pairwise pseudo-labels on $\mathcal{D}_{u}$, if the similarity is larger than a pre-defined threshold, then such pair of samples are recognized as similar pair. Then the model is trained on $\mathcal{D}_{u}$ with the binary pseudo-labels for novel category discovery. Such type of methods rely on an implicit assumption that general knowledge of what constitutes a good class is shared across old and new classes. AutoNovel [13] is a representative work in NCD, dividing NCD into three steps. The model is firstly pre-trained on $\mathcal{D}_{l}$ and $\mathcal{D}_{u}$ collectively with self-supervision on rotation-prediction [53] to grasp fundamental representation. They proposed a novel pseudo-labeling criteria namely ranking statistics to obtain pairwise similarities. Concretely, two samples are deemed to be similar if the top-k ranked dimensions of feature magnitudes are identical, which is robust and stable for knowledge transfer from old to novel classes. In the last step, the model is jointly optimized on $\mathcal{D}_{l}$ with classification loss and $\mathcal{D}_{u}$ with pseudo-labels, together with consistency regularization [93] for better representation learning. Some other methods advance NCD from different perspectives, including local part-level information [94], and neighborhood [95] in the feature space containing more potential positive samples. The adjacent samples in the feature space are valuable for constructing pseudo-positives and hard negatives for substantially better representation learning.
One-stage Methods. More recent works learn $\mathcal{D}_{l}$ and $\mathcal{D}_{u}$ simultaneously in a joint optimization manner. Zhong et al. [96] proposed to model old and novel classes into a joint label distribution, and leverage the mixup of labeled and unlabeled samples to generate more credible data for learning. UNO [97] is a simple yet effective method with a unified objective for NCD. UNO employs a swapped prediction task like SwAV [98] for pseudo-labeling followed by Sinkhorn-Knopp algorithm [99] to avoid trivial solutions. Kullback–Leibler divergence terms [100] to maximize the inter-class divergence and minimize the intra-class divergence could be incorporated to boost the performance.
@@
Figure 4: Illustrations of two types of methods of GCD, i.e., (A) non-parametric methods with K-means clustering and (B) parametric methods with learnable classification heads (bottom).
4.2 Generalized Category Discovery

### v3

@@ III-C Open-Set Recognition
The ability of a classifier to reject unknowns has been verified as highly correlated with its closed-set classification accuracy [258]. As a result, when applying some training tricks to the classifier for better closed-set classification, e.g., longer training, RandAugment [285], learning rate warmup, and label smoothing [286], the classifier consequently performs better in OSR. The authors [258] further found that applying baseline methods like MSP [10] and MLS [22] on a well-trained closed-set classifier could achieve competitively with or even outperform state-of-the-art OSR methods. Cen et al. [287] further revealed that wrong-classified ID samples have a similar uncertainty distribution to unknown samples, which limits the performance of OSR, and pre-training could enhance both the separation between ID and OOD samples and the separation between the correctly- and wrongly-classified ID samples.
2) Generative Models. Generative models provide different perspectives from discriminative models and could be utilized mainly in two ways, i.e., one is generative and probabilistic modeling of known and unknown classes, and the other is to explicitly generate fake unknown samples for greater separation between known and unknown classes. From the perspective of generative modeling, the basic idea is to model posterior distributions in the latent space to approximate conditional Gaussian distributions [288], and the probability of test samples could be used for unknown rejection. However, directly modeling the distribution is difficult, and most existing methods employ off-the-shelf generative models to generate pseudo-unknown samples to augment classifiers. For example, G-OpenMax [46] generates unknown samples with GANs [289] and fine-tunes the model with OpenMax [41] loss. OpenGAN [47] adopts an adversarial training framework to generate pseudo-unknown samples on the deep feature space. By further taking the difficulty [290] of the generated unknown samples into consideration, models could grasp fine-grained concepts of unknowns for results. Besides, without using explicit generative models, pseudo-unknown samples could be generated by optimizing a latent vector to have a small distance to training samples but produce low confidence on known categories [291]. These samples could be viewed as difficult open-space samples that are similar to known samples. A $(K+1)$-way classifier is trained on the combination of training samples and the generated counterfactual samples.
3) Hybrid Models. Discriminative and generative models are complementary to each other, and both have their specific advantages. In principle, discriminative models perform better on classification tasks than generative models. On the other hand, generative models could formalize the distribution of known and unknown samples, and additionally generate fake samples to augment discriminative models. As a consequence, combining both discriminative and generative models into hybrid models results in enormous potential for performing OSR tasks. Zhang et al. [48] proposed to train a flow-based generative model [292] and an inlier classifier on a shared feature space of the encoder. The flow-based model and the classifier are targeted to detect outliers and classify inliers, respectively, without interfering with each other. The flow-based model is trained on the feature space using the maximum likelihood loss function, which avoids the issue of assigning OOD with high probability [293] when learned in the input pixel space. Yang et al. [294, 49] creatively proposed convolutional prototype networks (CPN) for OSR with learnable class-wise prototypes serving as a classifier, which is different from conventional linear classifiers. The prototypical classifier could be viewed as Gaussian generative modeling for each class, and leaves more room for open set samples in the feature space. CPN is trained on both discriminative and generative loss functions to encourage inter-class separation and intra-class compactness, respectively, resulting in less open set risk and remarkable OSR performance. ARPL [295, 51] takes a similar spirit of CPN [49], but models class-wise reciprocal points representing the extra-class space, i.e., modeling the open space of each category. ARPL [51] additionally employs adversarial enhancement to generate diverse confusing samples to further reduce the open space risk caused by confusing samples. The class-wise points of CPN [49] and ARPL [51] could be generalized to class-wise continuous manifolds [52], which are constructed by class-wise auto-encoders upon the feature space. Each class-wise manifold could be viewed as an infinite number of prototype/reciprocal points, and thus more flexible and generalizable.
III-D Theoretical Study of Unknown Rejection
Despite numerous empirical approaches for unknown rejection, theoretical investigations remain limited. Regarding the difficulty of unknown rejection, existing works predominantly resort to openness [255] $O^{\star}$ defined by the number of known and unknown classes:
$O^{\star}=1-\sqrt{\frac{2\times K_{\text{ID}}}{K_{\text{ID}}+K_{\text{OOD}}}},$ (3)
where $K_{\text{ID}},K_{\text{OOD}}$ denote the number of known and unknown classes, respectively. The more unknown categories compared to the number of known categories, the more challenging the unknown rejection task becomes. Vaze et al. [258] proposed a more rational metric, namely semantic similarity [296]. Open-set samples that share a greater number of visual features with known training samples, such as attributes, color, and shape, are more difficult to reject and detect. Fang et al. [297] regarded OOD detection as a $(K+1)$-class classification task where $K$ is the number of known class, and then established a theoretical framework for the probably approximately correct learnability of OOD detection. This study finds a necessary condition for the learnability of OOD detection, and presents the impossibility theorems when in-distribution and out-of-distribution overlap with each other. Cheng et al. [30] explored whether and how expanding the training set using generated data can improve unknown rejection performance and offered several new insights. The authors established theoretical results that increasing ID generated training data can largely benefit OOD detection under the case where OOD inputs are roughly equally distant from all classes, while having little impact for case where OOD inputs are substantially closer to one class relative to the others. Zhu et al. [298, 25] revealed an interesting reliable overfitting phenomenon, i.e., the unknown rejection performance can be easily overfitting during the training of a model. Then, Cheng et al. [27] theoretically show that LASSO [299] largely eliminates the reliable overfitting based on a linear model.
There are also some works focusing on the relationship between OOD detection and misclassification detection. Specifically, it has been observed that many unknown rejection methods excel at detecting OOD samples but sacrifice the ability to filter out misclassified in-distribution examples [33, 300, 298], leaving models vulnerable in practical scenarios. To understand this phenomenon, Zhu et al. [298] demonstrated the misalignment between the rejection rules of misclassified ID samples and OOD samples by investigating the Bayes-optimal reject rules. Narasimhan et al. [301, 302] provided a theoretical analysis of the trade-off between OOD detection and misclassification detection.
III-E Evaluation of Unknown Rejection
The most widely used metric for OOD detection is the Area Under the Receiver Operating Characteristic curve [303] (AUROC), which is a threshold-independent metric that can be interpreted as the probability that a positive example (ID) is assigned a higher prediction score than a negative example. In addition, there are many other metrics like FPR95, AUPR [10] to imply how the ID and OOD samples are separated. Specifically, AUPR is the Area under the precision–recall (PR) curve. The PR curve is a graph showing the precision versus recall. AUPR typically regards the ID samples as positive. FPR95 can be interpreted as the probability that a negative example is misclassified as a correct prediction when the true positive rate (TPR) is as high as $95\%$. The true positive rate can be computed by TPR=TP/(TP+FN), where TP and FN denote the number of true positives and false negatives, respectively. We provide a comprehensive performance comparison of OOD detection and OSR in Table II and Table III. Specifically, for OOD detection, the ID datasets are CIFAR-10 and CIFAR-100, and the results are the average on the six common benchmarks [30, 298]: Textures [304], SVHN [305], Place365 [306], LSUN-C, LSUN-R [307] and iSUN [308]. From the results in Table II, we can observe that outlier-aided methods like OE [32] perform the best, which is reasonable since OE uses extra natural outlier images. Training-stage methods such as LogitNorm [24] and GenData [30] outperform various post-hoc approaches. For OSR, the evaluations are conducted on five datasets, including MNIST [309], SVHN [310], CIFAR [311] and TinyImageNet [312]. Recent hybrid methods like ARPL [51] and CSSR [52] achieve relatively remarkable performance, considering that they combine the advantages of both generative and discriminative models.
TABLE II: Comparative results of several OOD methods on CIFAR-10/100, and the results are the average on the six common benchmarks. Here, bold and underline indicate the best and the second-best results.
Method Venue ID dataset: CIFAR-10 ID dataset: CIFAR-100
AUROC↑ AUPR↑ FPR95↓ AUROC↑ AUPR↑ FPR95↓
Model: ResNet-18
MSP [10] ICLR 2017 92.40 98.21 47.51 80.88 95.45 74.46
Energy [12] NeurIPS 2020 94.32 98.49 28.78 78.21 94.62 79.89
MaxLogit [22] ICML 2022 94.25 98.52 29.89 78.03 94.58 79.97
ViM [14] CVPR 2022 94.41 98.75 29.62 80.57 95.44 74.52
KNN [16] ICML 2022 94.09 98.77 38.52 78.47 94.68 74.98
LogitNorm [24] ICML 2022 94.63 98.87 31.44 81.70 95.37 64.79
RegMixup [313] NeurIPS 2022 93.03 98.42 39.84 78.22 94.78 81.73
OE [32] ICLR 2019 98.61 99.71 4.78 89.70 97.72 52.98
MSP++ [258] ICLR 2022 93.21 98.36 41.36 77.03 94.36 80.38
GenData [30] IJCV 2024 96.27 99.21 21.32 84.24 96.39 66.07
Model: WRN-28-10
MSP [10] ICLR 2017 91.11 97.34 41.76 77.80 94.40 79.48
Energy [12] NeurIPS 2020 92.46 97.48 29.19 76.87 93.91 79.54
MaxLogit [22] ICML 2022 92.43 97.50 30.02 76.67 93.90 79.71
ViM [14] CVPR 2022 96.34 99.11 17.07 84.36 96.14 62.42
KNN [16] ICML 2022 94.73 98.91 33.38 79.34 94.95 75.89
LogitNorm [24] ICML 2022 97.36 99.42 13.26 85.27 96.50 64.61
RegMixup [313] NeurIPS 2022 93.16 98.29 36.92 83.13 95.93 69.38
OE [32] ICLR 2019 98.65 99.74 3.69 91.43 98.18 49.11
MSP++ [258] ICLR 2022 91.83 97.45 36.16 76.42 93.90 80.00
GenData [30] IJCV2024 96.83 99.21 15.18 84.93 96.45 64.76
TABLE III: Comparative results of several OSR methods with AUROC.
Method Venue MNIST SVHN CIFAR10 CIFAR+10 CIFAR+50 TinyImageNet
MSP [10] Baseline 97.8 88.6 67.7 81.6 80.5 57.7
OpenMax [41] CVPR 2016 98.1 89.4 69.5 81.7 79.6 57.6
G-OpenMax [46] BMVC 2017 98.4 89.6 67.5 82.7 81.9 58.0
OSRCI [291] ECCV 2018 98.8 91.0 69.9 83.8 82.7 58.6
CSOSR [284] CVPR 2019 99.1 89.9 88.3 91.2 90.5 58.9
C2AE [43] CVPR 2019 98.9 92.2 89.5 95.5 93.7 74.8
PROSER [42] CVPR 2021 96.4 94.3 89.1 96.0 95.3 69.3
OpenHybrid [48] ECCV 2020 99.5 94.7 95.0 96.2 95.5 79.3
DIAS [290] ECCV 2022 99.2 94.3 85.0 92.0 91.6 73.1
CPN [49] TPAMI 2020 99.0 92.6 82.8 88.1 87.9 63.9
ARPL [51] TPAMI 2021 99.6 96.3 90.1 96.5 94.3 76.2
ARPL+CS [51] TPAMI 2021 99.7 96.7 91.0 97.1 95.1 78.2
CSSR [52] TPAMI 2022 - 97.9 91.3 96.3 96.2 82.3
IV Novel Class Discovery
IV-A Problem Formulation and Key Challenges
@@
1) Problem Formulation. Novel class discovery [58, 314] is the second step in open-world machine learning, which aims to automatically discover novel categories from unlabeled data based on the model’s prior knowledge. Novel class discovery is an extension of unknown rejection, requiring models to not only reject unknown samples, but also further extrapolate to classify the rejected samples. Given a dataset $\mathcal{D}_{\text{train}}=\mathcal{D}_{l}\cup\mathcal{D}_{u}$ with two subsets, where $\mathcal{D}_{l}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N_{l}}\subset\mathcal{X}\times\mathcal{Y}_{l}$ is the labeled dataset and $\mathcal{D}_{u}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N_{u}}\subset\mathcal{X}\times\mathcal{Y}_{u}$ is the unlabeled data. It is worth noting that $\mathcal{Y}_{u}\neq\mathcal{Y}_{l}$, i.e., there are novel classes in the unlabeled data outside the old classes in $\mathcal{D}_{l}$. Let $\mathcal{C}_{\text{old}},\mathcal{C}_{\text{new}}$ denote the old and new classes in the labeled and unlabeled dataset, with the number of classes $K_{\text{old}}=|\mathcal{C}_{\text{old}}|,K_{\text{new}}=|\mathcal{C}_{\text{new}}|$, respectively. The objective is to cluster and discover $\mathcal{C}_{\text{new}}$ in $\mathcal{D}_{u}$ leveraging the knowledge learned on $\mathcal{D}_{l}$.
@@
In essence, novel class discovery is a clustering task on $\mathcal{D}_{u}$. Regarding its differences from unsupervised clustering [315, 316], the latter aims to cluster unlabeled data in a purely unsupervised manner without any prior knowledge of the classification criterion. As a result, unsupervised clustering is not a fully learnable task. For example, in the absence of any external information, the model might face a dilemma, for example, being confused about whether to cluster red flowers and red cars together according to colors or to cluster red flowers and yellow flowers together according to species. From this perspective, some labeled data $\mathcal{D}_{l}$ is necessary and could provide some prior knowledge indicating what constitutes a class to clarify the clustering criterion and eliminate ambiguity. In a word, novel class discovery is a more reasonable and pragmatic task, and could be referred to as deep transfer clustering [55], in which models are supposed to transfer the knowledge learned from old and labeled classes to cluster novel categories in the unlabeled data.
@@
2) Key Challenges. Intrinsically, the semantic similarity between new and old categories determines the extent of knowledge transferable from old to new categories, thereby influencing the performance of category discovery. If the semantic similarity is limited, the labeled old class might degrade the performance of novel class discovery [317]. Technically, novel category discovery could be referred to as open-world semi-supervised learning [80], where unlabeled samples contain new classes, which makes the conventional pseudo-labeling mechanism [318] not applicable due to the non-shared categories between labeled and unlabeled data. Moreover, in practical scenarios, the number of novel categories is always unknown, which makes the task challenging.
This section will sequentially introduce two primary tasks of novel class discovery, namely novel category discovery (Sec. IV-B) and generalized category discovery (Sec. IV-C), along with representative methods. We also elaborate on the class number estimation issue (Sec. IV-D), which is critical in realistic scenarios. Besides, we point out some extended settings of novel class discovery to reflect some new trends of this field (Sec. IV-E). Last, we summarize theoretical results of category discovery (Sec. IV-F) and then present the evaluation protocols and performance comparisons (Sec. IV-G).
Figure 6: The evolutionary tree of novel class discovery methods.
IV-B Novel Category Discovery
@@
1) Isolated-training Methods. Early works in NCD mainly transfer knowledge from $\mathcal{D}_{l}$ to $\mathcal{D}_{u}$ for category discovery in isolated steps. They first focus on $\mathcal{D}_{l}$ only to learn the similarity function. Subsequently, they utilize the learned similarity function to perform unlabeled learning exclusively on the unlabeled data $\mathcal{D}_{u}$. In other words, these methods do not learn new and old classes simultaneously. Instead, they learn them at separate stages. A pioneering work, DTC [55], formalizes NCD as a deep transfer clustering task, and builds the method upon unsupervised clustering DEC [315]. DTC first trains a model on $D_{l}$ in a supervised manner, which is then utilized to extract features of $\mathcal{D}_{u}$ to initialize cluster centroids. Then, DTC performs alternative optimization between the soft targets and model parameters, as well as the cluster centroids. Later works predominantly resort to pseudo-labels to transfer knowledge from labeled samples from old classes to unlabeled samples from novel classes. However, due to the disjoint classes, one could not simply adopt the naive pseudo-labels [318] in semi-supervised learning [319]. To overcome this issue, several methods rely on pairwise pseudo-labels [53, 54] indicating whether two samples are similar or not. Specifically, the model is first trained on $\mathcal{D}_{l}$ to assign pairwise pseudo-labels on $\mathcal{D}_{u}$, if the similarity is larger than a pre-defined threshold, then such pairs of samples are recognized as similar pairs. Then, the model is trained on $\mathcal{D}_{u}$ with the binary pseudo-labels for novel category discovery. Such types of methods rely on an implicit assumption that general knowledge of what constitutes a good class is shared across old and new classes. They assume that the similarity function learned from $\mathcal{D}_{l}$ is also applicable to $\mathcal{D}_{u}$. ResTune [56] disentangles the representations of old and new classes, which facilitates the adaptation of new classes and mitigates the forgetting of old classes.
@@
Figure 7: Illustrations of two types of methods of Generalized Category Discovery (GCD), i.e., (A) non-parametric methods with K-means clustering and (B) parametric methods with learnable classification heads.
2) Joint-training Methods. The fundamental principle of these methods lies in training the model simultaneously on both labeled $\mathcal{D}_{l}$ and unlabeled data $\mathcal{D}_{u}$. Although some of them [58, 62] employ multi-stage training, they utilize labeled data of old classes $\mathcal{D}_{l}$ during the clustering of unlabeled new classes in $\mathcal{D}_{u}$, rather than treating the learning of new and old classes as isolated, independent tasks, as in the isolated-training methods. AutoNovel [58] is a representative work in NCD, dividing NCD into three steps. The model is first pre-trained on $\mathcal{D}_{l}$ and $\mathcal{D}_{u}$ collectively with self-supervision on rotation prediction [267] to grasp fundamental representation. They proposed a novel pseudo-labeling criteria, namely ranking statistics to obtain pairwise similarities $s_{ij}$. Concretely, two samples are deemed similar if the top-k ranked dimensions of feature magnitudes are identical, which is robust and stable for knowledge transfer from old to novel classes, as in Eq. (4).
$s_{ij}=\mathds{1}(\text{top}_{k}\bm{z}_{i}^{u}=\text{top}_{k}\bm{z}_{j}^{u}),$ (4)
where $\bm{z}_{i}^{u}$ denotes the visual feature of the $i$-th unlabeled sample. In the last step, the model is jointly optimized on conventional cross-entropy loss on $\mathcal{D}_{l}$ and binary cross-entropy loss (Eq. (5)) on $\mathcal{D}_{u}$ with pair-wise pseudo-labels, together with consistency regularization [320] (Eq. (6)) for better representation learning.
$\mathcal{L}_{\text{bce}}=-\frac{1}{|\mathcal{B}|^{2}}\sum_{i\in\mathcal{B}}\sum_{j\in\mathcal{B}}\big{[}s_{ij}\log\eta^{u}(\bm{z}_{i}^{u})^{\top}\eta^{u}(\bm{z}_{j}^{u})$ (5)
$+(1-s_{ij})\log(1-\eta^{u}(\bm{z}_{i}^{u})^{\top}\eta^{u}(\bm{z}_{j}^{u}))\big{]},$
$\mathcal{L}_{\text{mse}}$ $=\frac{1}{|\mathcal{B}_{l}|}\sum_{i\in\mathcal{B}_{l}}\|\eta^{l}(\bm{z}_{i}^{l})-\eta^{l}(\hat{\bm{z}}_{i}^{l})\|^{2}$ (6)
$+\frac{1}{|\mathcal{B}_{u}|}\sum_{i\in\mathcal{B}_{u}}\|\eta^{l}(\bm{z}_{i}^{u})-\eta^{l}(\hat{\bm{z}}_{i}^{u})\|^{2},$
where $\eta^{l}(\cdot),\eta^{u}(\cdot)$ denotes the classification head for old and new classes, respectively, and $\bm{z}_{i}$ is the feature representation, while $\hat{\bm{z}}_{i}$ is another view with a different data augmentation. Some other methods advance NCD from different perspectives, including local part-level information [62], and neighborhood [59] in the feature space containing more potential positive samples. The adjacent samples in the feature space are valuable for constructing pseudo-positives and hard negatives for substantially better representation learning. Zhong et al. [60] model old and novel classes in a joint label distribution, and introduce OpenMix to MixUp [274] the labeled and unlabeled samples, as well as generating pseudo-labels by mixing respective pseudo and ground-truth labels, which are more credible for learning. UNO [61] is a simple yet effective method with a unified objective for NCD. UNO employs a swapped prediction task between two views motivated by SwAV [321]. The pseudo targets are derived through the Sinkhorn-Knopp algorithm [322], which encourages equipartition of pseudo-labels across various classes. The unlabeled self-learning and supervised learning on labeled data are collectively implemented in an end-to-end manner. ComEx [63] divides and conquers NCD with two groups of compositional experts, with each characterizing the whole dataset from complementary perspectives. Li et al. [64] propose to utilize Kullback–Leibler divergence terms to maximize the inter-class divergence and minimize the intra-class divergence, achieving strong performance.
IV-C Generalized Category Discovery